Question: "Each side of a square is 4 m long. When each side is increased by x m, the area is doubled. Find the value of x." Please help!
Answer by stanbon(75887)

"Each side of a square is 4 m long. When each side is increased by x m, the area is doubled. Find the value of x."
----
Original Area = 4^2 = 16 sq meters
Solve:
(4+x)^2 = 2*16
16 + 8x + x^2 = 32
x^2+8x-16 = 0
---
x = [-8 +- sqrt(64-4*1*-16)]/2
---
x = [-8 +- 8sqrt(2)]/2
---
Positive solution:
x = [-8 + 8sqrt(2)]/2
x = 1.6569
=================
Cheers,
Stan H.
=================